/the probability that a married man watches a certain TV show is 0,4and the probability taht a married mwoman watches the show is 0.5. The probability that a man watches the show agiven that his wife watches it is 0.7. Find the probabiltiyt that a married couple watches the show; a wife watches the show given that her husband does; and at least one person of a married c ouple will watch the show.


Answer by stanbon(29603)

the probability that a married man watches a certain TV show is 0.4
and the probability that a married woman watches the show is 0.5.
The probability that a man watches the show given that his wife watches is 0.7
---------------------------------
Find the probability that a married couple watches the show;
P(man AND woman) = P(man|woman)*P(woman) = 0.7*0.5 = 0.35
---------------------------------
a wife watches the show given that her husband does;
P(woman|man) = P(man AND woman)/P(man) = 0.35/0.4 = 0.8750
-----------------------------------
and at least one person of a married couple will watch the show.
P(at least 1) = 1 - P(neither)
--
Need to figure out P(neither watch):
P(both watch) = 0.35
P(man watches and woman does not) = 0.40-0.35 = 0.05
P(woman watches and man does not) = 0.50-0.35 = 0.15
-----------------------------------------
P(neither watches) = 0.35+0.15+0.05 = 0.55
--------------------------
Therefore P(at least 1 watches) = 1 - 0.55 = 0.45
=======================================================
Cheers,
Stan H.